# Clustering of timed sequences – Application to the analysis of care pathways

### v1

@@ 4.1 Average Timed Sequences
Let $\mathcal{S}={s_{1},\dots,s_{m}}$ be a set of timed sequences, Algorithm 2 details the steps for generating an average timed sequence inspired by the principle of vertical averages of DBA [14]. At line 1, the current average time sequence, $\mathcal{C}^{(0)}$ is initialized with a timed sequence chosen from $\mathcal{S}$. Then, the distance from $\mathcal{C}^{(0)}$ to each of the other sequences is calculated to retrieve the optimal alignment $M^{*}$ represented by the indices $i_{r}$ and $i_{s}$ of matching elements in each timed sequence. Lines 8 to 10 collect, on the one hand, the list of paired events in the sequences of $\mathcal{S}$ per event in $\mathcal{C}^{(0)}$ and, on the other hand, their dates per event in $\mathcal{C}^{(0)}$. These lists are then averaged lines 11-12. The average of events is constructed as the histogram of event distributions. Finally, line 13 creates a new average sequence which can be refined in a subsequent iteration until convergence. In our implementation, we assume convergence after a fixed number of iterations.
@@
The choice of the timed sequence in line 1 is crucial for the result. We choose to randomly select a sequence of the longest length. We will see that our algorithm decreases the length of the current average timed sequence, but it can not augment it.
Data : $\mathcal{S}=\{s_{1},s_{2},...,s_{m}\}$: list of timed sequences
@@ 2 $k\leftarrow 0$;
             $\mu,\pi^{k}_{i}\leftarrow DropDTW(\mathcal{C}^{(k)},s_{i})$ ;
              // calculate the drop-DTW and the best alignment between $\mathcal{C}^{(k)}$ and $s_{i}$
5             for $(i_{r},i_{s})\in\pi^{k}_{i}$ do
@@ 8       $k\leftarrow k+1$;
Figure 2 illustrates one iteration of average calculation for the three sequences of Example 1. It assumes that we already have a current average sequence $s_{r}$. The figure depicts the alignment by drop-DTW of $s_{r}$ with each sequence. $s^{\prime}_{r}$ is the resulting updated average sequence. As the $A$ event has been paired with both $A$ and $C$, the first probabilistic event of $s^{\prime}_{r}$ combines the two types of events (proportionally). As the first $B$ event of $s_{r}$ is paired with no event, it is removed from $s^{\prime}_{r}$. The other events are purely paired with similar event types, then the probabilistic event is a pure $C$ or $B$ in $s^{\prime}_{r}$ but their time positions have been updated.
@@ 4.2 Convergence of the Algorithm
The algorithm 2 refines iteratively an average sequence for a set of sequences, $\left(s_{i}\right)_{i\in[m]}$. In this algorithm, $\mathcal{C}^{(k)}$ denotes the current average sequence at the $k$-th iteration. At step $k$, the inertia of the average sequence is defined as:
@@
The inertia represents the average error incurred by using $\mathcal{C}^{(k)}$ as the representative of the set of sequences $\left(s_{i}\right)_{i\in[m]}$. In the following, we prove that $I^{(k+1)}$ is a decreasing suite, i.e., $I^{(k+1)}\leq I^{(k)}$, for all $k\in\mathbb{N}$.
@@
where $d$ is the metric between timed events (see Section 3.2) and $\pi$ plays the same role as the matrix $M$ introduced in section 3.1. It represents the alignment between two sequences. In the remainder, we denote by $\pi^{(k)}_{i}$ the optimal alignment of sequence $s_{i}$ with sequence $\mathcal{C}^{(k)}$, i.e., the alignment that minimizes the sum of costs.
@@
It is worth noticing that we retrieve the inertia of the $k+1$-th average sequence $I^{(k+1)}$ by summing the left-hand terms over $i$, and then:
@@
In the case $\left\{y\mid(x,y)\in\pi^{k}_{j}\right\}=\emptyset$, this means that for all $s_{i}$, there is an empty row at line $x$ of the alignment with $\mathcal{C}^{(k+1)}$. For instance, the first $B$ event in Figure 2 is assigned to none event. Then, this event contribute to the inertia only via a drop-cost. Removing such an event from $\mathcal{C}^{(k+1)}$ turns out to remove an empty row for all sequences $s_{j}$, and thus substract one to all $Pr(\pi^{k+1}_{j})$. Then, it also decreases the true values of $q_{j}^{(k+1)}$.
@@ 4.3 Parameters
The proposed method for averaging time sequences has two types of parameters: the parameters of the drop-DTW metric ($\tau\in\mathbb{R}$, $p_{e}\in[0,1]$, $p_{t}\in[0,1]$, $\delta\in\mathbb{R}$) and the parameter of the Algorithm 2 (maximum iterations for drop-DTW and iterative averaging).
@@
The ratio $\frac{p_{t}}{p_{e}}$ can be inferred from the delay $\tau$ by the following reasoning. For events located within the time interval $[t-\tau,t+\tau]$, the importance of the event type should be greater than the importance of time. We must then choose $p_{e}$ to be at least $\tau$ times greater than $p_{t}$. With lost of generality, the value of $p_{t}$ can be set at $1$ by default. The drop-cost is also derived from $\tau$, i.e., the delay from which there is no longer any interest in associating two events. We thus propose to define the drop-cost as:
$\delta=\frac{1}{\tau^{2}}(\tau+t_{max})^{2}+1$
@@
5 Clustering
The proposed metric between timed sequences and the averaging techniques are used to adapt two classical clustering techniques: a hierarchical ascendant clustering (HAC) and a K-means algorithm [10]. These two algorithms require to predefine the expected number of clusters, $K$ [^3] Model selection based on the inertia of the clustering may be applied to select the optimal number of clusters.
@@ 5 Clustering
The K-means also proceeds iteratively to refine $K$ clusters. K-means algorithm has two main steps: first the sequences of each cluster are averaged (algorithm 2). The computed averages are the representatives of their clusters. Then, new clusters are made by assigning optimally each sequence to one of the $K$ representative.
@@ 6 Experiments and Results on Synthetic Data
In this section, we present experiments on synthetic. We conducted several tests on synthetic data to evaluate the effectiveness of the method under various scenarios and to study the effect of the parameters.
For this first test, we created two model timed sequences that differ in one or more of the dimensions we want to study: sequence length, event nature, and event dates. Therefore, there are $8=2^{3}$ possible variations. A dataset consists of 15 sequences for each of the 8 possible variations and the base sequence (thus $135=9\times 15$ sequences in total). The dates of these timed sequences are obtained and made random by adding uniform noise to the date. The objective of this experiment is to evaluate the accuracy of identifying the 9 categories of sequences.
We tested two sets of parameters: first, we chose the ratio $\frac{p_{t}}{p_{e}}=\frac{1}{9}$, with $\delta=+\infty$. In this configuration, the cost of an event is more important than the time cost for 3 days around the event. The second set of parameters used was $\frac{p_{t}}{p_{e}}=\frac{1}{400}$, still with $\delta=+\infty$. This time, the difference cost between events is more important than the time cost over the first 20 days.
In the case of the first configuration, we observe that all clusters were correctly identified (Kappa score of $1$). However, in the case of the second configuration, we observed a degradation of results with a Kappa score of $0.62$. This degradation was expected because the maximum time distances observed between events are around 20. Consequently, the time notion is practically never taken into account by the metric.
@@ 6.1 Effect of Drop-Cost
We can then test the effect of the drop-cost on a synthetic dataset, with and without drop-cost. We take 3 model sequences, 2 of which are similar, except that one of them has an aberrant value.
@@
Intuitively, we want to cluster sequences 1 and 3 together. We apply HierAsTiSeq without drop-cost and obtain the confusion matrix in Table 1 on the left, with a Kappa score of $-0.5$ (disagreement). We observe that the algorithm associated the set from sequence 1 with the set from sequence 2.
Adding drop-cost coinciding with the distances between the sequences allows us to obtain the confusion matrix in Table 1 on the right, with a Kappa score of $1$ (total agreement). The drop-cost allows neglecting the last event considered aberrant, and the results match expectations.
@@
We then notice that all events dated from sequence 1 are present in sequence 3, but they are all different from sequence 2. We would therefore like to cluster sequences from 1 and 3 together. Keeping the same parameters as before. Without drop-cost, we obtain the confusion matrix in Table 2 on the left. This result corresponds to expectations because the sequences from 2 and 3 are then associated as they are closer in time. However, by adding a $\delta=4$, we obtain the confusion matrix in Table 2 on the right, which is logical because the events furthest in time are no longer associated.
The tests on synthetic values present the expected results, thus validating the effectiveness of clustering in the situations described. We now continue the study of the method with an experiment on real data to evaluate its ability to identify relevant clusters.
@@ 7 Application to Real-World Care Pathways
In this section, we apply the proposed method to real-world care pathways and compare our results with those of Sequence State Analysis (SSA) [6]. SSA, particularly the TraMineR tool, is commonly used for categorizing care pathways [17].
This experiment is conducted on the data warehouse of the Greater Paris Hospital (AP-HP). We identified a population of 3311 patients having a stay with a lung resection surgery. We collected events that occurred within 90 days around a lung resection surgery as well as the date of death if applicable. Events further away are clinically less interesting to study. For clinical reasons, we also decided to keep only the first occurrence of chemotherapy for each patient. We made the same choice for radiotherapy and immunotherapy. It is worth noting that deaths are excluded from the the care pathway. However, we keep this information for the survival analysis of the clusters.
The parameters chosen are $K=5$ clusters, $\tau=7$, $t_{\text{max}}=16$, and $\delta=1$. For choosing the number of clusters, different values were tested. By increasing the number of clusters, it is mainly small-sized clusters that split. Therefore, there may not necessarily be an interest in increasing this number. $\tau$ is set to 7 days because, in our dataset, it may be interesting to associate the same types of events that occurred within a week of each other. Finally, based on the dataset, we consider that there is no longer any interest in associating two events more than 16 days apart.
Refer to captionRefer to captionRefer to caption
Figure 3: Clusters obtained by the clustering methods, from left to right: HierAsTiSeq, K-means and TraMineR. Clusters are row-wised. Each cluster is represented by the histogram of event types over time. Each color corresponds to a specific type of event (see Figure 4 for color legend). The higher the bar, the more events there are at the given time across all patients. The vertical red line indicates the index-date of the resection (0 delay).
@@ 7.1 Comparison with TraMineR
Figure 3 corresponds to the 5 clusters for the three methods. Each cluster is represented as an histogram of events over time. One can first notice that TraMineR provides clusters that look very similar to each others. It is difficult to identify specificities for each cluster. Contrary to the other two methods, based on our the drop-DTW metric, the sizes of the clusters are more homogeneous. For HierAsTiSeq and K-means, there is one large cluster and several small clusters that gather patients specific timed events. The two large clusters (cluster 5 for HierAsTiSeq and cluster 1 for K-means) looks very similar in distribution of events. HierAsTiSeq identifies two clusters with specific behaviors: the cluster 1 has no post-operative events. This may correspond to patients that had the surgery in the hospital but their follow-up in another medical structure; the cluster 2 are patients with short delays between pre- or post-operative care events around the surgery. The clusters 2 and 3 identified by K-means also gather patients with a late pre-operative consultation (wrt the resection). According to the distributions, these patients have two pre-operative consultations: in cluster 2, the first pre-operative consultation occurs earlier than in cluster 3; and cluster 3 holds patients having a late post-operative consultation.
@@
It is worth noting that results of K-means with drop-DTW may significantly differ across different runs due to the randomness of the cluster seeds. HierAsTiSeq and TraMineR are more deterministic, but the iterative computation of drop-DTW or optimal matching may introduce some randomness. Nonetheless, the hierarchical clustering strategy of HierAsTiSeq and TraMineR is resource demanding. This prevents to apply them on the entire dataset. This was possible only with K-means and takes only a few minutes (on 1 core of the AP-HP servers).
Figure 4: Clusters of care-pathways identified by K-means with drop-DTW on 3211 patients. Each row is a cluster with the representation of the barycenter on the left and the representation of the events distribution over time on the right. Each color corresponds to a specific type of event.

### v2

@@ 4.1 Average Timed Sequences
Let $\mathcal{S}=\{s_{1},\dots,s_{m}\}$ be a set of timed sequences, Algorithm 2 details the steps for generating an average timed sequence inspired by the principle of vertical averages of DBA (Petitjean et al., 2011). In this algorithm, $\mathcal{C}^{(k)}$ denotes the current average sequence at the $k$-th iteration. At line 1, the current average time sequence, $\mathcal{C}^{(0)}$ is initialized with a timed sequence chosen from $\mathcal{S}$. Then, this current average is processed into the outer loop which repeats $maxit$ times the alignment and update steps. The alignment step (line 6-11) calculates the alignment by drop-DTW between $\mathcal{C}^{(0)}$ and each of the other sequences. The alignment to one sequence is obtained line 8 from the distance matrix computed the drop-DTW. We refer the reader to the original DTW algorithm to get the details on how to compute an alignment from a distance matrix. Lines 9 to 11 collect, on the one hand, the list of paired events in the sequences of $\mathcal{S}$ per event id, $e$, in $\mathcal{E}_{e}$ and, on the other hand, their dates per event id, $e$, in $\mathcal{T}_{e}$. These lists are then averaged lines 12-13. This is the update step. This vertical average of events is constructed as the histogram of event distributions. Lastly, line 14 creates a new average sequence which can be refined in a subsequent iteration until convergence. In our implementation, we assume convergence after a fixed number of iterations but a convergence criteria could be implemented to prevent useless iterations.
@@
The choice of the timed sequence in line 1 is crucial for the result. We choose to randomly select a sequence of the longest length. We made this choice because our algorithm decreases the length of the current average timed sequence, but it can not increase it. Note that there are some smarter initialization strategies for k-means like algorithms that could improve the convergence, for instance K-means++ (Arthur and Vassilvitskii, 2007), or recent provably good seedings methods (Bachem et al., 2016).
Data : $\mathcal{S}=\{s_{1},s_{2},...,s_{m}\}$: list of timed sequences; $maxit\in\mathbb{N}$: maximum number of iterations
@@ 2 $k\leftarrow 0$;
             $\mu,D\leftarrow DropDTW(\mathcal{C}^{(k)},s_{i})$ ;
              // calculate the distances matrix of drop-DTW between $\mathcal{C}^{(k)}$ and $s_{i}$
             $\pi^{*}\leftarrow GetAlignement(D)$ ;
              // calculate the best alignment from $D$
5             for $(i_{r},i_{s})\in\pi^{*}$ do
@@ 8       $k\leftarrow k+1$;
Example 4 (Iteration of the averaging algorithm).
Figure 2 illustrates one iteration of average calculation for the three sequences of Example 1. It assumes that we already have a current average sequence $s_{r}$ (a surgery, then a consultation and two radiotherapy sessions). The figure depicts the alignments by drop-DTW of $s_{r}$ with each sequence of Example 1. The gray lines represents the alignments (line 8). $s^{\prime}_{r}$ is the updated average sequence obtained by the lines 11-12 of Algorithm 2. As the $C$ event has been paired with both $C$ and $R$ events, the second probabilistic event of $s^{\prime}_{r}$ combines the two types of events (proportionally). In addition, the first $R$ event of $s_{r}$ is removed from $s^{\prime}_{r}$ because it is paired with no event. The other events are purely paired with similar event types, then the probabilistic event is a pure $S$ or $R$ in $s^{\prime}_{r}$ but their time positions have been adjusted with respect to the positions of the new aligned events.
@@ 4.2 Convergence of the Algorithm
In this section, we prove the theoretical convergence of the proposed algorithm.
The algorithm 2 refines iteratively an average sequence for a set of sequences, $\left\{s_{i}\right\}_{i\in[m]}$. At step $k$, the inertia of the average sequence is defined as:
@@
The inertia represents the average error incurred by using $\mathcal{C}^{(k)}$ as the representative of the set of sequences $\left\{s_{i}\right\}_{i\in[m]}$. To demonstrate the convergence, we prove that $I^{(k+1)}$ is a decreasing suite, i.e., $I^{(k+1)}\leq I^{(k)}$, for all $k\in\mathbb{N}$.
@@
where $d$ is the metric between timed events (see Section 3.2) and $\pi$ plays the same role as in section 3.1. It represents the alignment between two sequences. In the remainder, we denote by $\pi^{(k)}_{i}$ the optimal alignment of sequence $s_{i}$ with sequence $\mathcal{C}^{(k)}$, i.e., the alignment that minimizes the sum of costs.
@@
It is worth noticing that we retrieve the inertia of the $(k+1)$-th average sequence $I^{(k+1)}$ by summing the left-hand terms over $i$, and then:
@@
In the case $\left\{y\mid(x,y)\in\pi^{k}_{j}\right\}=\emptyset$, this means that for all $s_{i}$, there is an empty row at line $x$ of the alignment with $\mathcal{C}^{(k+1)}$. For instance, the first $R$ event in Figure 2 is assigned to none event. Then, this event contributes to the inertia only via a drop-cost. Removing such an event from $\mathcal{C}^{(k+1)}$ turns out to remove an empty row for all sequences $s_{j}$, and thus subtracts one to all $Pr(\pi^{k+1}_{j})$. Then, it also decreases the true values of $q_{j}^{(k+1)}$.
@@ 4.3 Parameters
The proposed method for averaging time sequences has two types of parameters: the parameters of the drop-DTW metric ($\tau\in\mathbb{R}$, $p_{e}\in[0,1]$, $p_{t}\in[0,1]$, $\delta\in\mathbb{R}$) and the parameters of the Algorithm 2 (maximum iterations for drop-DTW and iterative averaging).
@@
The ratio $\frac{p_{t}}{p_{e}}$ can be inferred from the delay $\tau$ by the following reasoning. For events located within the time interval $[t-\tau,t+\tau]$, the importance of the event type should be greater than the importance of time. We must then choose $p_{e}$ to be at least $\tau^{2}$ times greater than $p_{t}$. Without lost of generality, the value of $p_{t}$ can be set at $1$ by default. The drop-cost is also derived from $\tau$, i.e., the delay from which there is no longer any interest in associating two events. We thus propose to define the drop-cost as:
$\delta=\frac{1}{\tau^{2}}(\tau+t_{max})+1$
@@
Note that this suggested values are only a guideline to initiate a set up for the metric. Nonetheless, it has to be adjusted and tested on real data.
5 Clustering of Timed Sequences
The proposed metric between timed sequences and the averaging techniques are used to adapt two classical clustering techniques: a hierarchical ascendant clustering (HAC) and a K-means algorithm (Hartigan, 1975). These two algorithms require to predefine the expected number of clusters, $K$ [^3] Model selection based on the inertia of the clustering may be applied to select the optimal number of clusters.
@@ 5 Clustering of Timed Sequences
The $K$-means also proceeds iteratively to refine $K$ clusters. K-means algorithm has two main steps: first the sequences of each cluster are averaged (Algorithm 2). The computed averages are the representatives of their clusters. Then, new clusters are made by assigning optimally each sequence to one of the $K$ representatives.
@@ 6 Experiments and Results on Synthetic Data
In this section, we present experiments on synthetic data. We conducted several tests to evaluate the effectiveness of the method under various scenarios and to study the effect of the parameters.
6.1 Effects of the $\frac{p_{t}}{p_{e}}$ Ratio
For this first experiment, we created two model timed sequences that differ in one or more of the dimensions we want to study: sequence length, event nature, and event dates. Therefore, there are $8=2^{3}$ possible variations. A dataset consists of 15 sequences for each of the 8 possible variations and the base sequence (thus $135=9\times 15$ sequences in total). The dates of these timed sequences are obtained and made random by adding uniform noise to the date. The objective is to evaluate the accuracy of identifying the 9 categories of sequences.
We tested two sets of parameters: first, we chose the ratio $\frac{p_{t}}{p_{e}}=\frac{1}{9}$, with $\delta=+\infty$. With this configuration, the cost of an event is more important than the time cost for 3 days around the event. The second set of parameters used was $\frac{p_{t}}{p_{e}}=\frac{1}{400}$, still with $\delta=+\infty$. This time, the difference cost between events is more important than the time cost over the first 20 days.
In the case of the first configuration, we observe that all clusters generated by HierAsTiSeq were correctly identified (Kappa score of $1$). However, in the case of the second configuration, we observe a degradation of results with a Kappa score of $0.62$. This degradation was expected because the maximum time distance observed between events are around 20. Consequently, the notion of time is practically never taken into account by the metric.
@@ 6.2 Effect of Drop-Cost
We can then test the effect of the drop-cost on a synthetic dataset, with and without drop-cost. We consider 3 models of sequences. We generate 15 concrete sequences from each of them by randomizing the date of the events [^4] We added a centered normal random noise with a standard deviation of 1.. The sequences of the first and the third types are similar, except that the sequences of the third type have an aberrant event.
@@
Intuitively, we want to cluster together all sequences of types 1 and 3. We apply HierAsTiSeq without drop-cost and obtain the confusion matrix in Table 1 on the left, with a Kappa score of $-0.5$ (disagreement). We observe that the algorithm associated instances of sequence 1 with instances of sequence 2.
Adding drop-cost coinciding with the distances between the sequences allows us to obtain the confusion matrix in Table 1 on the right, with a Kappa score of $1$ (total agreement). The drop cost allows neglecting the last event considered aberrant, and the results match expectations.
@@
We then notice that all events dated from sequence 1 are present in sequence 3, but they are all different from sequence 2. We would therefore like to cluster sequences from 1 and 3 together. Keeping the same parameters as before. Without drop-cost, we obtain the confusion matrix in Table 2 on the left. This result corresponds to expectations because the sequences from 2 and 3 are then associated as they are closer in time. However, with $\delta=4$, we obtain the confusion matrix in Table 2. The table on the right, as expected, shows that the events furthest in time are no longer associated.
The tests on synthetic data lead to the expected results, thus validating the effectiveness of clustering in the situations described. We now continue studying the method with an experiment on real data to evaluate its ability to identify relevant clusters.
@@ 7 Application to Real-World Care Pathways
In this section, we apply the proposed method to real-world care pathways and compare our results with those of Sequence State Analysis (SSA) (Gabadinho et al., 2011). SSA, particularly the TraMineR tool, is commonly used for categorizing care pathways (Roux et al., 2019; Le Meur et al., 2019).
This experiment is conducted on the data warehouse of the Greater Paris Hospital (AP-HP). We identified a population of 3 31133113\,3113 311 patients having a stay with a lung resection surgery. We collected events that occurred within 90 days around a lung resection surgery as well as the date of death if applicable. Events further away are clinically less interesting to study. For clinical reasons, we also decided to keep only the first occurrence of chemotherapy for each patient. We made the same choice for radiotherapy and immunotherapy.
The parameters chosen are $K=5$ clusters, $\tau=7$, $t_{\text{max}}=16$, and $\delta=1$. For choosing the number of clusters, different values were tested. By increasing the number of clusters, it is mainly small-sized clusters that split. Therefore, there may not necessarily be an interest in increasing this number. $\tau$ is set to 7 days because, in our dataset, it may be interesting to associate the same types of events that occurred within a week of each other. Finally, based on the dataset, we consider that there is no interested in pairing two events more than 16 days apart.
For this experiment, we used the servers from the Paris University Hospital made of AMD EPYC 7713 processors. Our algorithms are neither parallelized nor distributed. However, the implementation of softDTW by Tavenard et al. (2020) can utilize multiple cores. During our experiments, we requested only one core from the servers.
Figure 3: Clusters obtained by the clustering methods, from left to right: HierAsTiSeq, K-means and TraMineR. Clusters are ordered row-wise by their size. Each cluster is represented by the histogram of event types over time. Each color corresponds to a specific type of event (see Figure 4 for color legend). The higher the bar, the more events there are at the given time across all patients. The vertical red line indicates the index-date of the resection (0 delay).
@@ 7.1 Comparison with TraMineR
Figure 3 corresponds to the $K=5$ clusters for the three methods. Each cluster is represented as an histogram of events over time. One can first notice that TraMineR provides clusters that look very similar to each others. It is difficult to identify specificities for each cluster. Contrary to the other two methods, based on our the drop-DTW metric, the sizes of the clusters are more homogeneous. For HierAsTiSeq and K-means, there is one large cluster and several small clusters that gather patients specific timed events. The two large clusters (cluster 1 for HierAsTiSeq and cluster 1 for K-means) looks very similar in distribution of events. HierAsTiSeq identifies two clusters with specific behaviors: the cluster 3 has no post-operative events. This may correspond to patients that had the surgery in the hospital but their follow-up in another medical structure; the cluster 2 are patients with short delays between pre- or post-operative care events around the surgery. The clusters 3 and 4 identified by K-means also gather patients with a late pre-operative consultation (wrt the resection). According to the distributions, these patients have two pre-operative consultations: in cluster 3, the first pre-operative consultation occurs earlier than in cluster 4; and cluster 4 holds patients having a late post-operative consultation.
@@
It is worth noting that results of K-means with drop-DTW may significantly differ across different runs due to the randomness of the cluster seeds. HierAsTiSeq and TraMineR are more deterministic, but the iterative computation of drop-DTW or optimal matching may introduce some randomness. Nonetheless, the hierarchical clustering strategy of HierAsTiSeq and TraMineR is resource demanding. This prevents to apply them on the entire dataset. This was possible only with K-means and takes only a few minutes.
Figure 4: Clusters of care-pathways identified by K-means with drop-DTW on 3 31133113\,3113 311 patients. Each row is a cluster with the representation of the barycenter on the left and the representation of the events distribution over time on the right. Each color corresponds to a specific type of event.